An electric field E=20i+30j exists in space.if the potential at the origin is taken to be zero,then what is the potential at (2m,2m)?

Grade:12

5 Answers

Arun
5 years ago
Dear student
 
Change in potential = - integral (vector E dot vector dl), dl is length element 
So V-0 = -vector E dot integral(vector dl) from (0,0) to (2,2) 

V= -(20i ̂+30j ̂ ) . {(2,2)-(0,0)} 
V= -(20,30) . (2,2) 
V= -(40+60) 
V= -100V
Khimraj
5 years ago
potential change due to x component of E in y direction is zero as angle is 90 and similialy
potential change due to y component of E in x direction is zero
SO change in potential
V – 0 = – ( 20*2 + 30 *2) = -100V.
Keshav
4 years ago
V=-ED 
V=(20i+30j)*(2-0,2-0)
V=(20*2+30*2)
V=(-1)(100)
V=-100V 
Change in potential = - integral (vector E dot vector dl), dl is length element 
So V-0 = -vector E dot integral(vector dl) from (0,0) to (2,2) 
 
Yash Chourasiya
3 years ago
Hello Student

ΔV = −∫E.dl
where,dlisthelengthelement
V−0 = −E.∫dl {Limit from (0,0) to(2,2) }
V = −(20iˆ+30jˆ).{(2,2)−(0,0)}
V = −(20iˆ+30jˆ).(2iˆ+2jˆ)
V = −(40+60)
V = −100V

I hope this answer will help you.
ankit singh
3 years ago
dear student ΔV = −∫E.dlwhere,dlisthelengthelementV−0 = −E.∫dl {Limit from (0,0) to(2,2) }V = −(20iˆ+30jˆ).{(2,2)−(0,0)}V = −(20iˆ+30jˆ).(2iˆ+2jˆ)V = −(40+60)V = −100V